Situations that can have probabilities can involve data or unsure outcomes that can be used to commute into chances of failure or successes. For example, using contingency tables, we could compute simple probabilities of successes with data given. For example,  

So for this class lets say we have the students, 10 female and 10 male. You can use this contingency chart to calculate the probability of male students that may fail for example or vice versa. Using the table below, calculate the percentage of male and female students likely to pass the class.

Pass Fail

Male 7 3

Female 8 2
Answer #1

Here n(m) = 10 , n(f) = 10

So, p(male and female students likely to pass the class)

= 8/20 + 7/20

= 15/20 = 0.75

So, 75% is the answer of this